# Weak interaction axial form factors of the octet baryons in nuclear medium

### v1

@@ VI.5 Summary
In the present section we exemplify how to use the calculations of the electromagnetic and axial-vector form factors in-medium to calculate the single-differential cross sections of NC and CC reactions in-medium. In general the cross sections are reduced in-medium. The magnitude of the calculations depend on the incident particle (neutrino or antineutrino). The calculation of the $\frac{d\sigma}{dQ^{2}}$ in terms of $Q^{2}$ requires the knowledge of the neutrino/antineutrino energy $E_{\nu}$ distribution.
The present calculations are consistent with the dominance of the neutrino cross sections over the antineutrino single-differential cross sections. The effect was observed on the BNL experiments [83, 85] where the averaged neutrino over the antineutrino single-differential cross sections for muon neutrinos ($\nu_{\mu}p\to\nu_{\mu}p$ vs $\bar{\nu}_{\mu}p\to\bar{\nu}_{\mu}p$) are determined. The main difference to the present calculations are on the lepton family (muon neutrinos instead of electron neutrinos) and the energy range of neutrinos ($E_{\nu}\simeq 1$ GeV, instead of $E_{\nu}\simeq 0.1$ GeV).
@@ VII Outlook and conclusions
The literature about the electromagnetic and axial structure of baryons in the nuclear medium is scarce, except for the case of the nucleon. This work is one more contribution to the field. We present a systematic study of the axial form factors of the octet baryons in terms of the square transfer momentum ($q^{2}=-Q^{2}$) and the density of the nuclear medium. We extend here previous studies in the free space for octet baryons immersed in a nuclear medium. The calculations presented can be used in future for the studies of interactions between the octet baryons with neutrinos or antineutrinos in the nuclear medium, at low, intermediate and large $Q^{2}$. We expect that the present study stimulates future works on the subject and promote a debate between the different formalisms used in the studies of the reactions in the nuclear medium.
Our calculations are based on the covariant spectator quark model, which takes into account the valence quark effects and the meson cloud dressing effects on the electroweak interactions with baryon systems. The model has been used successfully in the studies of the electromagnetic and axial structure of several baryon systems including resonances of the nucleon in free space. The extension of the model to the nuclear medium takes into account the modifications of the properties of mesons and baryons (masses and coupling constants) in the medium determined by the quark-meson coupling model.
@@
Our calculations in the nuclear medium suggest that the weak axial-vector form factor $G_{A}$ is suppressed in the nuclear medium. The suppression is stronger for light baryons and high densities. The effect increases in general with $Q^{2}$. In the case of the $n\to p$ transition about 30% for the nuclear density for $Q^{2}=2$ GeV2.
As for the induced pseudoscalar form factor $G_{P}$, because it is an observable that is hard to measure even in the vacuum, we have made more qualitative estimates. Our calculations suggest that $G_{P}$ is significantly suppressed in medium. In the case of $\Delta I=1$ transitions (only $u$ and $d$ quarks are involved) the suppression is more significant for light baryons and for transitions with significant suppression in $G_{A}$. In the case of $\Delta S=1$ transitions (a strange quark transition is involved), one can have significant suppression at low $Q^{2}$ in general. At large $Q^{2}$ the suppression can be relevant for light baryons and very mild for heavy baryons (especially for $\Xi$).
@@ Acknowledgements.
G. R. thanks H. Parada for useful discussions. G. R. was supported by the Basic Science Research Program funded by the Republic of Korea Ministry of Education (Grant No. NRF–2021R1A6A1A03043957). K.T. was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil), Processes No. 313063/2018-4 and No. 426150/2018-0, and FAPESP Process No. 2019/00763-0 and No. 2023/07313-6, and his work was also part of the projects, Instituto Nacional de Ciência e Tecnologia - Nuclear Physics and Applications (INCT-FNA), Brazil, Process No. 464898/2014-5. The work of M. K. C. is supported by the National Research Foundation of Korea (Grant Nos. NRF–2021R1A6A1A03043957 and NRF–2020R1A2C3006177).
@@ Appendix A Gell-Mann matrices
For the discussion of the quark electromagnetic current it is convenient to define

### v2

@@ VI.5 Summary
In the present section, we exemplify how to use the calculations of the electromagnetic and axial-vector form factors in medium to calculate the single-differential cross sections of NC and CC reactions in medium. In general the cross sections are reduced in medium. The magnitude of the calculations depends on the incident particle (neutrino or antineutrino). The calculation of the $\frac{d\sigma}{dQ^{2}}$ in terms of $Q^{2}$ requires the knowledge of the neutrino/antineutrino energy $E_{\nu}$ distribution.
The present calculations are consistent with the dominance of the neutrino cross sections over the antineutrino single-differential cross sections. The effect was observed on the BNL experiments [128, 130], where the averaged neutrino over the antineutrino single-differential cross sections for muon neutrinos ($\nu_{\mu}p\to\nu_{\mu}p$ vs $\bar{\nu}_{\mu}p\to\bar{\nu}_{\mu}p$) are determined. The main difference to the present calculations is on the lepton family (muon neutrinos instead of electron neutrinos) and the energy range of neutrinos ($E_{\nu}\simeq 1$ GeV, instead of $E_{\nu}\simeq 0.1$ GeV).
@@ VII Outlook and conclusions
The literature about the electromagnetic and axial structure of baryons in the nuclear medium is scarce, except for the case of the nucleon. This work is one more relevant contribution to the field. We present a systematic study of the axial form factors of the octet baryons in terms of the square transfer momentum ($q^{2}=-Q^{2}$) and the density of the nuclear medium. We extend here previous studies in the free space for octet baryons immersed in a nuclear medium. The calculations presented can be used in the future studies of interactions between the octet baryons with neutrinos or antineutrinos in the nuclear medium, at low, intermediate and large $Q^{2}$. We expect the present study to stimulate future works on the subject and promote a debate between the different formalisms used in the studies of the reactions in the nuclear medium.
Our calculations are based on the covariant spectator quark model, which takes into account the valence quark effects and the meson cloud dressing effects on the electroweak interactions with baryon systems. The model has been used successfully in studies of the electromagnetic and axial structure of several baryon systems including resonances of the nucleon in free space. The extension of the model to the nuclear medium takes into account the modifications of the properties of mesons and baryons (masses and coupling constants) in the medium determined by the quark-meson coupling model.
@@
Our calculations in the nuclear medium suggest that the weak axial-vector form factor $G_{A}$ is suppressed in the nuclear medium. The suppression is stronger for light baryons and high densities. The effect increases, in general, with $Q^{2}$. In the case of the $n\to p$ transition about 30% for the nuclear density for $Q^{2}=2$ GeV2.
As for the induced pseudoscalar form factor $G_{P}$, because it is an observable that is hard to measure even in the vacuum, we have made more qualitative estimates. Our calculations suggest that $G_{P}$ is significantly suppressed in medium. In the case of $|\Delta I|=1$ transitions (only $u$ and $d$ quarks are involved) the suppression is more significant for light baryons and for transitions with significant suppression in $G_{A}$. In the case of $|\Delta S|=1$ transitions (a strange quark transition is involved), one can have significant suppression at low $Q^{2}$ in general. At large $Q^{2}$ the suppression can be relevant for light baryons and very mild for heavy baryons (especially for $\Xi$).
@@ Acknowledgements.
G. R. thanks H. Parada for useful discussions. G. R. was supported by the Basic Science Research Program funded by the Republic of Korea Ministry of Education (Grant No. NRF–2021R1A6A1A03043957). K.T. was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil), Processes No. 313063/2018-4, No. 426150/2018-0 and No. 3014199/2022-2, and FAPESP Processes No. 2019/00763-0 and No. 2023/07313-6, and his work was also part of the projects, Instituto Nacional de Ciência e Tecnologia - Nuclear Physics and Applications (INCT-FNA), Brazil, Process No. 464898/2014-5. The work of M. -K. C. is supported by the National Research Foundation of Korea (Grants Nos. NRF–2021R1A6A1A03043957 and NRF–2020R1A2C3006177).
@@ Appendix A Gell-Mann matrices
For the discussion of the quark electromagnetic current, it is convenient to define
@@
Appendix B Numerical results for the axial-vector form factor $G_{A}$
We present here the numerical results for the form factors $G_{A}$ for all transitions between octet baryon members discussed in Secs. V.1, V.2 and V.3. For a more clear comparison of magnitudes and shapes, we present the results of equivalent functions in the same column.
In the cases where the final result $G_{A}$ is negative, we represent the function $-G_{A}$, for an easier comparison with the other transitions.
In Fig. B1 we present the results for the $|\Delta I|=1$ transitions, and in Figs. B2 and B3 we present the results for the $|\Delta S|=1$ transitions.
The axial-vector form factors in the vacuum are on the right column [Sec. V.1]. The contributions from the quark core are represented by the dashed lines, and the contributions from the meson cloud by the dashed-doted lines. The final results correspond to the solid lines. On the left column we present the final results of $G_{A}$ for the vacuum ($\rho=0$) and medium (densities $\rho=0.5\rho_{0}$ and $\rho=\rho_{0}$) [Sec. V.2].
In Figs. B4 and B5 we represent the ratios $G_{A}^{\ast}(Q^{2})/G_{A}(Q^{2})$ on the right column, and the ratios $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$ on the left panel, for the $|\Delta I|=1$ and $|\Delta S|=1$ cases, respectively [Sec. V.3].
Concerning the $|\Delta I|=1$ transitions (Figs. B1 and B4), we recall that the form factors $\Sigma^{+}\to\Lambda$ and $\Sigma^{-}\to\Lambda$ differ by a sign, and that the form factors $\Sigma^{-}\to\Sigma^{0}$ and $\Sigma^{0}\to\Sigma^{+}$ are equal.
As for the $|\Delta S|=1$ transitions (Figs. B2, B3 and B5), the ratios for the transitions $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$ are the same. The same is true also for the $\Xi^{0}\to\Sigma^{+}$ and $\Xi^{-}\to\Sigma^{0}$ transitions.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure B1: Axial-vector form factor $G_{A}$ for $|\Delta I|=1$ transitions. Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure B2: Axial-vector form factor $G_{A}$ for $|\Delta S|=1$ transitions (part 1). Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure B3: Axial-vector form factor $G_{A}$ for $|\Delta S|=1$ transitions (part 2). Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure B4: Axial-vector form factor $G_{A}$ for $|\Delta I|=1$ transitions. Ratios $G_{A}^{\ast}/G_{A}$ and $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure B5: Axial-vector form factor $G_{A}$ for $|\Delta S|=1$ transitions. Ratios $G_{A}^{\ast}/G_{A}$ and $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$. The ratios are the same for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$. The ratios are the same for $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$.
Appendix C Numerical results for the induced pseudoscalar form factor $G_{P}$
We present here the numerical results for the form factors $G_{P}$ for all transitions between octet baryon members discussed in Secs. V.1, V.2 and V.3. For a more clear comparison of magnitudes and shapes we present the results of equivalent functions in the same column.
As for the function $G_{A}$, in the cases where the final result $G_{P}$ is negative, we represent the function $-G_{P}$.
In Fig. C1 we present the results for the $|\Delta I|=1$ transitions and in Figs. C2 and C3 we present the results for the $|\Delta S|=1$ transitions.
On the right column we present the induced pseudoscalar form factors in the vacuum [Sec. V.1]. In the figures, we discriminate the result for the bare contribution, the combination of the bare and pseudoscalar meson pole contribution, and the sum of all the contributions (Bare + Pole + Meson Cloud). On the left column we present the final results for the vacuum ($\rho=0$) and medium (densities $\rho=0.5\rho_{0}$ and $\rho=\rho_{0}$) [Sec. V.2].
In Figs. C4 and C5, we represent the ratios $G_{P}^{\ast}(Q^{2})/G_{P}(Q^{2})$ on the right column, and and the ratio $G_{P}^{\ast}(Q^{2})/G_{P}(Q_{p}^{2})$ where $Q_{p}^{2}=0.012$ GeV2, on the left column, for the $|\Delta I|=1$ and $|\Delta S|=1$ cases, respectively [Sec. V.3].
As in the case of the form factor $G_{A}$, there are some relations between the theoretical calculations associated with different channels. For the $|\Delta I|=1$ transitions (Figs. C1 and C4), we recall that the form factors $\Sigma^{+}\to\Lambda$ and $\Sigma^{-}\to\Lambda$ differ by a sign, and that the form factors $\Sigma^{-}\to\Sigma^{0}$ and $\Sigma^{0}\to\Sigma^{+}$ are the same.
As for the $|\Delta S|=1$ transitions (Figs. C2, C3 and C5), the ratios for the transitions $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$ are the same. The same property is also valid for the $\Xi^{0}\to\Sigma^{+}$ and $\Xi^{-}\to\Sigma^{0}$ transitions.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure C1: Induced pseudoscalar form factor $G_{P}$ for $|\Delta I|=1$ transitions. Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure C2: Induced pseudoscalar form factor $G_{P}$ for $|\Delta S|=1$ transitions (part 1). Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure C3: Induced pseudoscalar form factor $G_{P}$ for $|\Delta S|=1$ transitions (part 2). Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure C4: Induced pseudoscalar form factor $G_{P}$ for $|\Delta I|=1$ transitions. Ratios $G_{P}^{\ast}/G_{P}$ and $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure C5: Induced pseudoscalar form factor $G_{P}$ for $|\Delta S|=1$ transitions. Ratios $G_{P}^{\ast}/G_{P}$ and $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$. The ratios are the same for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$. The ratios are the same for $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$.